Problem • Assume that it is May 15, 2010, and the U.S. Treasury has just issued securities with a May 2015 maturity, $1,000 par value, and a 2.2% coupon rate with semiannual coupons. • Since the origi...

Problem

Assume that it is May 15, 2010, and the U.S. Treasury has

just issued securities with a May 2015 maturity, $1,000 par

value, and a 2.2% coupon rate with semiannual coupons.

Since the original maturity is only five years, these would

be called “notes” as opposed to “bonds.”

The first coupon payment will be paid on November 15,

2010.

What cash flows will you receive if you hold this note until

maturity?   

##I need to know how to solve this problem with BAll Plus calculator.##
Answer #1

(Nov 2010) Cash flow 1 = 2.2%*1000/2 = 11

(May 2011) CF 2= 11

(Nov 2011) CF 3 = 11

(May 2012) CF 4 = 11

(Nov 2012) CF 5 = 11

(May 2013) CF 6 = 11

(Nov 2013) CF 7 = 11

(May 2014) CF 8 = 11

(Nov 2014) CF 9 = 11

(May 2015) CF 10= 11 + 1000 = 1011